1. If 5.00g of cucumber slices, contain 4.689g of water, what is the mass/mass percent of water

in the cucumber slices?

Mass/mass percent____________%

2. If the watch glass weighs 4.50g and the watch glass + food slices weigh 5.23g, what is the

mass of the food slices?

Mass of food slices_________g

3. If 9.00 moles of NaCl are dissolved to make a 3.50L aqueous solution of sodium chloride,

what is the Molarity of the solution?

_______________M NaCl solutio
Answer #1

1.

mass of cucumber slices   = 5g

mass of water                      = 4.689g

total mass of cucumber slices + water = 5 + 4.689 =   9.689g

mass percent of cucumber slices = mass of slice*100/total mass of cucumber +water

                                                       = 5*100/9.689 = 51.6%

2.

The watch glass weighs    = 4.50g

the watch glass + food slices weigh = 5.23g

the mass of the food slices   = the watch glass + food slices weigh -The watch glass weighs

                                          = 5.23-4.5   = 0.73g

The mass of the food slices = 0.73g

3.

no of moles of NaCl = 9 moles

volume of solution = 3.50L

molarity of NaCl   = no of moles/volume of solution in L

                           = 9/3.5

                            = 2.57M NaCl solution